OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1)

OpenStudy (andu1854):

first find the slope of the two points: -1 - 1/5-4 = -2/1 = -2 slope... Now pick either point and set it up like this: y -1 = -2 (x-4) and then solve for y to get y = mx + b form, with m = -2 as your slope